Circular No. 7641

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001ck IN UGC 9425 W. D. Li, University of California at Berkeley, reports the discovery

by LOTOSS (cf. IAUC 7514) of an apparent supernova in unfiltered images taken with the Katzman Automatic Imaging Telescope (KAIT) on June 3.3 (mag ∼ 17.2) and 4.3 UT (mag ∼ 16.8). The new object is located at α = 14^{h}37^{m}50^{s}.08, δ = +30^{o}29′02′′.2 (equinox 2000.0), which is 5′′.8 west and 3′′.0 north of the northern nucleus of UGC 9425. A KAIT image taken on May 24.3 showed nothing at this position (limiting mag ∼ 18.5).

SUPERNOVA 2001cl IN NGC 7260 Li also reports the LOTOSS discovery by M. Papenkova of another

apparent supernova (mag ∼ 17.2) in unfiltered images taken with the KAIT on June 4.5 and 5.4 UT. The new object is located at α = 22^{h}22^{m}37^{s}.52, δ = −4^{o}07′30′′.8 (equinox 2000.0), which is 15′′.5 east and 16′′.0 south of the nucleus of NGC 7260. A KAIT image of the same field taken on 2000 Dec. 5.2 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2001cm IN NGC 5965 X. J. Jiang and Y. L. Qiu, on behalf of Beijing Astronomical Obser-

vatory supernova survey, report the discovery of an apparent supernova found on unfiltered images taken on June 3.62 (mag ∼ 17.5) and 5.72 UT (mag ∼ 16.8). SN 2001cm is located at α = 15^{h}34^{m}06^{s}.70, δ = +56^{o}41′26′′.2 (equinox 2000.0), which is 36′′.2 east and 16′′.4 north of the nucleus of NGC 5965. An image taken on June 2.60 shows nothing at this position (limiting mag ∼ 18.0).

COMETS C/2001 K6, C/2001 K7, AND C/2001 K8 (SOHO) Further to IAUC 7634, D. Hammer reports his measurements for three

additional Kreutz sungrazing comets found by M. Oates, K. Černis, and T. Scarmato on LASCO C2 coronagraph images posted at the SOHO website. The reduced astrometry and orbital elements by B. G. Marsden appear on MPEC 2001-L09.

Comet  2001 UT  α_{2000}  δ_{2000}
C/2001 K6 May 23.368  4^{h}01.^{m}8  +18^{o}58′
C/2001 K7  23.493  4 02.0  +19 00
C/2001 K8  29.504  4 25.2  +19 54
2001 June 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green